OpenStudy (anonymous):

What is the solution to the equation log2 x - log2 4 =5?

OpenStudy (anonymous):

u mean log 2x and log 24?

OpenStudy (anonymous):

The rules for logarithms is: n^logn(x) = x

OpenStudy (anonymous):

Is the base 2??

OpenStudy (anonymous):

\[\log_{2} x = \log_{2} 4 =5\]

OpenStudy (anonymous):

* -

OpenStudy (anonymous):

rule* So, when your base is 2 (log2), then 2^log2(x)=x

OpenStudy (anonymous):

\[\log _{2} x - \log_{2} 4 = 5\] sorry that is the equation

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

See if you can eliminate the logarithms from the equation :)

OpenStudy (anonymous):

would it be x - 4 =5?

OpenStudy (anonymous):

Not quite, when you use an operation on one side of the equation, you must do it on the other side, otherwise you change the equation.

OpenStudy (anonymous):

You are close though :) 2^log2(4)=4 indeed

OpenStudy (anonymous):

hm.. what would I do to the other side of the equation??

OpenStudy (anonymous):

What did you do to the left side?

OpenStudy (anonymous):

I eliminated the log part

OpenStudy (anonymous):

How is log2(x) different from x

OpenStudy (anonymous):

log2 (x) - log2(4) = 5 log2 (x/4) = 5 2^5 = x/4 Finish her up.

OpenStudy (anonymous):

How did 2^5 come out? I'm confused..

OpenStudy (anonymous):

log(base) (answer) = (exponent) 2 is the base, 5 is the exponent, x/4 is the answer.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

but why would that become 2^5?

OpenStudy (anonymous):

where did the log go?

OpenStudy (anonymous):

Log is just another way to write out an exponent. So 2^5 is the same thing as log2 (x) = 5.

OpenStudy (anonymous):

For the last step he did: \[2^{log2(\frac{x}{4})}=2^5\]

OpenStudy (anonymous):

I see..

OpenStudy (anonymous):

Okay, I get it now.. so all I have to do now is to solve for x from 2^5 = x/4 right?

OpenStudy (anonymous):

Yep. You got it!

OpenStudy (anonymous):

o.o yay thanks!

OpenStudy (anonymous):

One more thing-- if there's two logarithms being subtracted do I always just divide them if they have the same base?

OpenStudy (anonymous):

Yes. And if they are added with the same base it's multiplication instead.

OpenStudy (anonymous):

then if they're multiplied it would be addition? and subtraction if they are divided?

OpenStudy (anonymous):

No, I don't remember exactly how it's done off of the top of my head but if the expanded forms are multiplied or divided then you should be able to multuply them together to get your answer. Let me see if I can remember how it's done.

OpenStudy (anonymous):

Okay, thanks :)

OpenStudy (anonymous):

I guess I should clarify. If you have something like log(1000/100) then it would be equal to log(1000) - log(100). However, if you have log(1000)/log(100) it is not.

OpenStudy (anonymous):

ohh.. i see

OpenStudy (anonymous):

That makes a lot sense. Thank you!

OpenStudy (anonymous):

Not a problem! =) I can't remember any shortcut for solving the latter but chances are you won't need a shortcut, just solve each log independently then find the difference of their answers.

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (unklerhaukus):

\[\boxed{\log_b x+\log_by=\log_bxy}\]\[\boxed{\log_b x-\log_by=\log_b\frac xy}\]

OpenStudy (anonymous):

That cleans it up, thank you :)